40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 57 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  14251   Sat Oct 13 20:11:10 2018 yukiUpdateComputer Scripts / Programsloss measurements
Quote:

the script "armloss_AS_calc.py",

  • "ETM_YAW_L_DEMOD_I_OUTPUT" was little noisy even when the arm was locked.
  • The reflected beam power when locked was higher than when misaligned. It seemed strange for me at first. Johannes suggested that it was caused by over-coupling cavity. It is possible when r_{ETMY}>>r1_{ITMY}.

Some changes were made in the script for getting the signals of beam power:

  • The script sees "C1:ASS-X(Y)ARM_ETM_PIT/YAW_L_DEMOD_I_OUTPUT" and stops running until the signals become small, however some offset could be on the signal. So I changed it into waiting until (DEMOD - OFFSET) becomes small. (Yesterday I wrote ETM_YAW_L_DEMOD_I_OUTPUT was about 15 and was little noisy. I was wrong. That was just a offset value.)
  • I added a code which stops running the script when the power of transmitted IR beam is low. You can set this threshold. The nominal value of "C1:LSC-TRX(Y)_OUT16" is 1.2 (1.0), so the threshold is set 0.8 now.  

In the yesterday measurement the beam power of ASDC is higher when locked than when misaligned and I wrote it maybe caused by over-coupled cavity. Then I did a calculation as following to explain this:

  • assume power transmissivity of ITM and ETM are 1.4e-2 and 1.4e-5.
  • assume loss-less mirror, you can calculate amplitude reflectivity of ITM and ETM.
  • consider a cavity which consists two mirrors and is loss-less, then \frac{E_{r}}{E_{in}} = \frac{-r_1+r_2e^{i\phi}}{1-r_1r_2e^{i\phi}} holds. r1 and r2 are amplitude reflectivity of ITM and ETM, and E is electric filed.
  • Then you can calculate the power of reflected beam when resonated and when anti-resonated. The fraction of these value is \frac{P_{RESONANT}}{P_{ANTI-RESO}} = 0.996, which is smaller than 1.
  • I found this calculation was wrong! Above calculatation only holds when cavity is aligned, not when misaligned. 99.04% of incident beam power reflects when locked, and (100-1.4)% reflects when misaligned. The proportion is P(locked)/P(misaligned)=1.004, higher than 1.

 

  14248   Fri Oct 12 20:20:29 2018 yukiUpdateComputer Scripts / Programsloss measurements

I ran the script for measuring arm-loss and calculated rough Y-arm round trip loss temporally. The result was 89.6ppm. (The error should be considered later.)

The measurement was done as follows:

  1. install hardware
    1. Put a PD (PDA520) in anti-symmetric beam path to get ASDC signal.
    2. Use a PD (PDA255) in MC2 box to get tranmitted light of IMC. It is used to normalize the beam power.
    3. Connect those 2 PDs to oscilloscope (IP: 192.168.113.25) and insert an internet cable to it.
  2. measure DARK noise
    1. Block beam going into PDs with dampers and turn off the room light.
    2. Run the script "armloss_dcrefl_acdcpd_scope.py" using "DARK" mode.
  3. measure the ASDC power when Y-arm locked and misaligned
    1. Remove dampers and turn off the room light.
    2. Dither-align the interferometer with both arms locked. Freeze outputs when done. (Click C1ASS.adl>!MoreScripts>ON and click C1ASS.adl>!MoreScripts>FreezeOutputs.)
    3. Misalign ETMX + ITMX. (Just click "Misalign" button.)
    4. Further misalign ITMX with the slider. (see previous study: ITMX needs to be misaligned further. Moving the slider by at least +0.2 is plentiful to not have the other beam interfere with the measurement.)
    5. Start the script "armloss_dcrefl_acdcpd_scope.py" using "ETMY" mode, which does the following:
      1. Resume dithering of the YARM.
      2. Check YARM dither error signal rms with CDS. If they're calm enough, proceed. (In the previous study the rms threshold was 0.7. Now "ETM_YAW_L_DEMOD_I" signal was 15 (noisy), then the threshold was set 17.)
      3. Freeze dithering.
      4. Start a new set of averages on the scope, wait T_WAIT (5 seconds).
      5. Read data (= ASDC power and MC2 trans) from scope and save.
      6. Misalign ETMY and wait 5s. (I added a code which switchs LSC mode ON and OFF.)
      7. Read data from scope and save.
      8. Repeat desired amount of times.
  4. calculate the arm loss
    1. Start the script "armloss_AS_calc.py", whose content is follows:
      • requires given parameters: Mode-Matching effeciency, modulation depth, transmissivity. I used the same value as Johannes did last year, which are (huga)
      • reads datafile of beam power at ASDC and MC2 trans, which file is created by "armloss_dcrefl_acdcpd_scope.py".
      • calculates arm loss from the equation (see 12528 and 12854).

Result:

YARM
('AS_DARK =', '0.0019517200000000003') #dark noise at ASDC 
('MC_DARK =', '0.02792') #dark noise at MC2 trans
('AS_LOCKED =', '2.04293') #beam power at ASDC when the cavity was locked 
('MC_LOCKED =', '2.6951620000000003')
('AS_MISALIGNED =', '2.0445439999999997') #beam power at ASDC when the cavity was misaligned
('MC_MISALIGNED =', '2.665312')

\hat{P} = \frac{P_{AS}-P_{AS}^{DARK}}{P_{MC}-P_{MC}^{DARK}} #normalized beam power 

\hat{P}^{LOCKED}=0.765,\ \hat{P}^{MISALIGNED}=0.775,\ \mathcal{L}=89.6\ \mathrm{ppm}

Comments:

  • "ETM_YAW_L_DEMOD_I_OUTPUT" was little noisy even when the arm was locked.
  • The reflected beam power when locked was higher than when misaligned. It seemed strange for me at first. Johannes suggested that it was caused by over-coupling cavity. It is possible when r_{ETMY}>>r1_{ITMY}.
  • My first (wrong) measurement said the arm loss was negative(!). That was caused by lack of enough misalignment of another arm mirrors. If you don't misalign ITMX enough then the beam or scattered light from X-arm would bring bad. The calculated negative loss would be appeared only when \frac{\hat{P}^{LOCKED}}{\hat{P}^{MISALIGNED}} > 1 + T_{ITM}
  • Error should be considered.
  • Parameters given this time should be measured again. 
  14247   Fri Oct 12 17:37:03 2018 SteveUpdateVACpressure gauge choices

We want to measure the pressure gradient in the 40m IFO

Our old MKS cold cathodes are out of order. The existing working gauge at the pumpspool is InstruTech CCM501

The plan is to purchase 3 new gauges for ETMY, BS and MC2 location.

Basic cold cathode     or    Bayard-Alpert Pirani

    

 

  14245   Fri Oct 12 12:29:34 2018 yukiUpdateComputer Scripts / Programsloss measurements

With Gautam's help, Y-arm was locked. Then I ran the script "armloss_dcrefl_asdcpd_scope.py" which gets the signals from oscilloscope. It ran and got data, but I found some problems.

  1. It seemed that a process which makes arm cavity mislaigned in the script didn't work.
  2. The script "armloss_dcrefl_asdcpd_scope.py" gets the signal and the another script "armloss_AS_calc.py" calculates the arm loss. But output file the former makes doesn't match with the type the latter requires. A script converts format is needed.

Anyway, I got the data needed so I will calculate the loss after converting the format.

  14244   Fri Oct 12 08:27:05 2018 SteveUpdateVACdrypump

Gautam and Steve,

Our TP3 drypump seal is at 360 mT [0.25A load on small turbo]  after one year.  We tried to swap in old spare drypump with new tip  seal. It was blowing it's fuse, so we could not do it.

Noisy aux drypump turned on and opened to TP3 foreline [ two drypumps are in  the foreline now ]  The pressure is 48 mT and 0.17A load on small turbo.

Attachment 1: forepump.png
forepump.png
  14243   Thu Oct 11 13:40:51 2018 yukiUpdateComputer Scripts / Programsloss measurements
Quote:

This is the procedure I follow when I take these measurements for the XARM (symmetric under XARM <-> YARM):

  1. Dither-align the interferometer with both arms locked. Freeze outputs when done.
  2. Misalign ETMY + ITMY.
  3. ITMY needs to be misaligned further. Moving the slider by at least +0.2 is plentiful to not have the other beam interfere with the measurement.
  4. Start the script, which does the following:
    1. Resume dithering of the XARM
    2. Check XARM dither error signal rms with CDS. If they're calm enough, proceed.
    3. Freeze dithering
    4. Start a new set of averages on the scope, wait T_WAIT (5 seconds)
    5. Read data (= ASDC power and MC2 trans) from scope and save
    6. Misalign ETMX and wait 5s
    7. Read data from scope and save
    8. Repeat desired amount of times
  5. Close the PSL shutter and measure the PD dark levels

Information for the armloss measurement:

  • Script which gets the data:  /users/johannes/40m/armloss/scripts/armloss_scope/armloss_dcrefl_asdcpd_scope.py
  • Script which calculates the loss: /users/johannes/40m/armloss/scripts/misc/armloss_AS_calc.py
  • Before doing the procedure Johannes wrote you have to prepare as follows:
    • put a PD in anti-symmetric beam path to get ASDC signal.
    • put a PD in MC2 box to get tranmitted light of IMC. It is used to normalize the beam power.
    • connect those 2 PDs to oscilloscope and insert an internet cable to it.
  • Usage: python2 armloss_dcrefl_asdcpd_scope.py [IP address of Scope] [ScopeCH for AS] [ScopeCH for MC] [Num of iteration] [ArmMode]

Note: The scripts uses httplib2 module. You have to install it if you don't have.

The locked arms are needed to calculate armloss but the alignment of PMC is deadly bad now. So at first I will make it aligned. (Gautam aligned it and PMC is locked now.) 

gautam: The PMC alignment was fine, the problem was that the c1psl slow machine had become unresponsive, which prevented the PMC length servo from functioning correctly. I rebooted the machine and undid the alignment changes Yuki had made on the PSL table.

  14241   Wed Oct 10 12:38:27 2018 yukiConfigurationLSCAll hardware was installed

I connected DAC - AIboard - PZTdriver - PZT mirrors and made sure the PZT mirrors were moving when changing the signal from DAC. Tomorrow I will prepare alignment servo with green beam for Y-arm.

  14240   Tue Oct 9 23:03:43 2018 yukiConfigurationLSCYarm Green locking was recovered

[ Yuki, Gautam, Steve ]

To align the green beam in Y-end these hardware were installed:

  • PZT mirrors in Y-end table
  • PZT driver in 1Y4 rack
  • Anti-Imaging board in 1Y4 rack
  • cables (DAC - AIboard - PZTdriver - PZT)
  • high voltage supplier 

I made sure that DAC CH9~16 and cable to AI-board worked correctly. 

When we applied +100V to PZT driver and connected DAC, AI-board and PZT drive, the output voltage of the driver was not correct. I'll check it tomorrow.

Attachment 1: Pic_1Y4.jpg
Pic_1Y4.jpg
Attachment 2: Pic_PZTcable.jpg
Pic_PZTcable.jpg
  14239   Tue Oct 9 16:05:29 2018 gautamConfigurationASCc1tst deleted, c1asy deployed.

Setting up c1asy:

  • Backed up old c1tst.mdl as c1tst_old_bak.mdl in /opt/rtcds/userapps/release/cds/c1/models
  • Copied the c1tst model to /opt/rtcds/userapps/release/isc/c1/models/c1asy.mdl as this is where the c1asx.mdl file resides.
  • Backed up original c1rfm.mdl as c1rfm_old.mdl in /opt/rtcds/userapps/release/cds/c1/models (since the old c1tst had an RFM block which is unnecessary).
  • Deleted offending RFM block from c1rfm.mdl.
  • Recompiled and re-installed c1rfm.mdl. Model has not yet been restarted, as I'd like suspension watchdogs to be shutdown, but c1susaux EPICS channels are presently not responsive.
  • Removed c1tst model (C-node91) from /opt/rtcds/caltech/c1/target/gds/param/testpoints.
  • Removed /opt/rtcds/caltech/c1/target/gds/param/tpchn_c1tst.par (at this point, DCUID 91 is free for use by c1asy).
  • Moved c1tst line in /opt/rtcds/caltech/c1/target/daqd/master to "old model definitions models" section.
  • Added /opt/rtcds/caltech/c1/target/gds/param/tpchn_c1asy.par to the master file.
  • Edited/diskless/root.jessie/etc/rtsystab to allow c1asy to be run on c1iscey.
  • Finally, I followed the instructions here to get the channels into frames and make all the indicators green.

Now Yuki can work on copying the simulink model (copy c1asx structure) and implementing the autoalignment servo.

Attachment 1: CDSoverview_ASY.png
CDSoverview_ASY.png
  14238   Mon Oct 8 18:56:52 2018 gautamConfigurationASCc1asx filter coefficient file missing

While pointing Yuki to the c1asx servo system, I noticed that the filter file for c1asx is missing in the usual chans directory. Why? Backups for it exist in the filter_archive subdirectory. But there is no current file. Clearly this doesn't seems to affect the realtime code execution as the ASX model seems to run just fine. I copied the latest backup version from the archive area into the chans directory for now.

  14237   Mon Oct 8 00:46:35 2018 yukiConfigurationASCY-end table upgrade
Quote:

I measured it with the wrong setting of a powermeter. The correct ones are here:

  • beam power going into doubling crystal (red12): 240 mW, 1064nm
  • beam power transmitted dichroic mirror (Y5): 0.70 mW, 532nm
  • beam power going into faraday isolator (green5): 0.58 mW, 532nm
  • beam power going out from faraday isolator (green5): 0.54 mW, 532nm
  • beam power going to ETMY: 0.37 mW, 532nm
  • beam power of transmitted green light of Y-arm, which was measured by C1:ALS-TRY_OUT: 0.5 (see attachment #1)

After installation I measured these power again.

  • beam power going into doubling crystal: 241 mW, 1064nm
  • beam power transmitted dichroic mirror: 0.70 mW, 532nm
  • beam power going into faraday isolator: 0.56 mW, 532nm
  • beam power going out from faraday isolator: 0.53 mW, 532nm
  • beam power going to ETMY: 0.36 mW, 532nm

There is a little power loss. That may be because of adding one lens in the beam path. I think it is allowable margin.

  14236   Sun Oct 7 22:30:42 2018 yukiConfigurationLSCYarm Green locking was recovered

I finished installation of optics in the Y-end and recovered green locking. Current ALS-TRY_OUTPUT is about 0.25, which is lower than before. So I still continue the alignment of the beam. The simulation code was attached. (Sorry. The optic shown as QWP2 is NOT QWP. It's HWP.)

Attachment 1: Pic_NewLayout1007.jpg
Pic_NewLayout1007.jpg
Attachment 2: YendGreenModeMatching.zip
  14235   Sun Oct 7 16:51:03 2018 gautamConfigurationLSCYarm triggering changed

To facilitate Yuki's alignment of the EY green beam into the Yarm cavity, I have changed the LSC triggering and PowNorm settings to use only the reflected light from the cavity to do the locking of Arm Cavity length to PSL. Running the configure script should restore the usual TRY triggering settings. Also, the X arm optics were macroscopically misaligned in order to be able to lock in this configuration.

  14234   Fri Oct 5 22:49:22 2018 yukiConfigurationASCY-end table upgrade

I designed a new layout. It has good mode-matching efficiency, reasonable sensitivity to component positions, good Gouy phase separation. I'm setting optics in the Y-end table. The layout will be optimized again after finishing (rough) installation.  (The picture will be posted later)

  14233   Fri Oct 5 17:47:55 2018 gautamConfigurationASCY-end table upgrade

What about just copying the Xend layout? I think it has good MM (per calculations), reasonable (in)sensitivity to component positions, good Gouy phase separation, and I think it is good to have the same layout at both ends. Since the green waist has the same size and location in the doubling crystal, it should be possible to adapt the X end solution to the Yend table pretty easily I think.

Quote:

The setup I designed is here. It can bring 100% mode-matching and good separation of degrees of TEM01, however I found a probrem. The picture of setup is attached #3. You can see the reflection angle at Y7 and Y8 is not appropriate. I will consider the schematic again.

  14232   Fri Oct 5 17:32:38 2018 yukiConfigurationASCY-end table upgrade

I measured it with the wrong setting of a powermeter. The correct ones are here:

  • beam power going into doubling crystal (red12): 240 mW, 1064nm
  • beam power transmitted dichroic mirror (Y5): 0.70 mW, 532nm
  • beam power going into faraday isolator (green5): 0.58 mW, 532nm
  • beam power going out from faraday isolator (green5): 0.54 mW, 532nm
  • beam power going to ETMY: 0.37 mW, 532nm
  • beam power of transmitted green light of Y-arm, which was measured by C1:ALS-TRY_OUT: 0.5 (see attachment #1)

The calculated conversion efficiency of SHG crystal is 1.2%W.

  14231   Fri Oct 5 00:46:17 2018 KojiConfigurationASCY-end table upgrade

???

The SHG crystal has the conversion efficiency of ~2%W (i.e. if you have 1W input @1064, you get 2% conversion efficiency ->20mW@532nm)

It is not possible to produce 0.58mW@532nm from 20.9mW@1064nm because this is already 2.8% efficiency.

 

  14230   Thu Oct 4 22:15:30 2018 yukiConfigurationASCY-end table upgrade

Before changing setup at Y-end table, I measured the status value of the former setup as follows. These values will be compared to those of upgraded setup.

  • beam power going into doubling crystal (red12): 20.9 mW with filter, 1064nm
  • beam power going out from doubling crystal (red12): 26.7 mW with filter, 532nm
  • beam power going into faraday isolator (green5): 0.58 mW without filter, 532nm
  • beam power going out from faraday isolator (green5): 0.54 mW without filter, 532nm
  • beam power going to ETMY: 0.37 mW without filter, 532nm
  • beam power of transmitted green light of Y-arm, which was measured by C1:ALS-TRY_OUT: 0.5 (see attachment #1)

(These numbers are shown in the attachment #2.)

The setup I designed is here. It can bring 100% mode-matching and good separation of degrees of TEM01, however I found a probrem. The picture of setup is attached #3. You can see the reflection angle at Y7 and Y8 is not appropriate. I will consider the schematic again.

Attachment 1: AUXYGreenLock20180921.jpg
AUXYGreenLock20180921.jpg
Attachment 2: Pic_FormerSetup.jpeg
Pic_FormerSetup.jpeg
Attachment 3: Pic_CurrentSetup1004.jpg
Pic_CurrentSetup1004.jpg
  14229   Thu Oct 4 08:25:50 2018 SteveUpdateVACrga scan pd81 at day 78

 

 

Attachment 1: pd81d78.png
pd81d78.png
  14228   Thu Oct 4 00:44:50 2018 yukiConfigurationASCAI board improvement

[ Yuki, Gautam ]

I made a cable which connects DAC port (40 pins) and AI board (25 pins). I will check if it works.

Tomorrow I will change setup for improvement of AUX Y-end green locking. Any optics for IR will not be moved in my design, so this work doesn't affect Y-arm locking with main beam. 
While doing this work, I will do:

  • check if the cable works
  • make another cable which connects AI board (10 pins) and PZT driver (10 pins).
  • check if eurocate in Y-rack (IY4?) applies +/-5V, +/-15V and +/-24V. It will be done using an expansion card.
  • improve alignment servo for X-end.
  • setup alignment servo for Y-end.
  • about optical loss measurement.  
  14227   Wed Oct 3 18:15:34 2018 yukiConfigurationASCAI board improvement

[ Yuki, Gautam ]

I improved Anti-Imaging board (D000186-Rev.D), which will be put between DAC port and PZT driver board.

It had notches at f = 16.6 kHz and 32.7 kHz, you can see them in the plot attached. So I replaced some resistors as follows:

  • R6 and R7 replaced with 511 ohm (1206 thin film resistor)
  • R8 replaced with 255 ohm (1206 thin film resistor)
  • R14 and R15 replaced with 549 ohm (1206 thin film resistor)
  • R16 replaced with 274 ohm (1206 thin film resistor)

Then the notch moved to 65.9 kHz (> sampling frequency of DAC = 64 kHz, good!). 
(The plot enlarged around the notch frequency and the plot of all channels will be posted later.)

All electronics and optics seem to be ready. 

Reference, elog:40m/8857
Diagram, D000186-D.pdf

Attachment 1: TF_AIboard.pdf
TF_AIboard.pdf
  14226   Wed Oct 3 14:24:40 2018 yukiConfigurationASCY end table upgrade plan

Interim Procedure Report:

Purpose

The current setup of AUX Y-arm Green locking has to be improved because:

  • current efficiency of mode matching is about 50%
  • current setup doesn't separate the degrees of freedom of TEM01 with PZT mirrors (the difference of gouy phase between PZT mirrors should be around 90 deg) 
  • we want to remotely control PZT mirrors for alignment

What to do

  • Design the new setup and order optices needed (finished!)
      - As the new setup I designed, adding a new lens and slightly changing the position of optics are only needed. The new lens was arrived here.
  • Check electronics (PZT, PZT driver, high voltage, cable, anti-imaging board) (finished!)
      
    - All electronics were made sure performing well.
      - The left thing to do is making a cable. (Today's tasks)
  • Calibrate PZT mirror [mrad/V] (finished!)
      
    - The result was posted here --> elog:40m/14224.
  • Measure the status value of the current setup (power of transmitted light ...etc) (Tomorrow, --> finished!)
  • Install them in the Y-end table and align the beam (Will start from Tomorrow) (The setup has a probrem I found on 10/04)
  • Measure the status value of the new setup
      - I want to finish above during my stay.
  • Prepare the code of making alignment automaticaly
  14225   Tue Oct 2 23:57:16 2018 gautamUpdatePonderSqueezeSqueezing scenarios

[kevin, gautam]

We have been working on double checking the noise budget calculations. We wanted to evaluate the amount of squeezing for a few different scenarios that vary in cost and time. Here are the findings:

Squeezing scenarios

Sqz [dBvac] fmin [Hz] PPRM [W] PBS [W] TPRM [%] TSRM [%]
-0.41 215 0.8 40 5.637 9.903
-0.58 230 1.7 80 5.637 9.903
-1.05 250 1.7 150 1 17
-2.26 340 10 900 1 17

All calculations done with

  • 4.5kohm series resistance on ETMs, 15kohms on ITMs, 25kohm on slow path on all four TMs.
  • Detuning of SRC = -0.01 deg.
  • Homodyne angle = 89.5 deg.
  • Homodyne QE = 0.9. 
  • Arm losses is 20ppm RT.
  • LO beam assumed to be extracted from PR2 transmission, and is ~20ppm of circulating power in PRC.

Scenarios:

  1. Existing setup, new RC folding mirrors for PRG of ~45.
  2. Existing setup, send Innolight (Edwin) for repair (= diode replacement?) and hope we get 1.7 W on back of PRM.
  3. Repair Innolight, new PRM and SRM, former for higher PRG, latter for higher DARM pole.
  4. Same as #3, but with 10 W input power on back of PRM (i.e. assuming we get a fiber amp).

Remarks:

  • The errors on the small dB numbers is large - 1% change in model parameters (e.g. arm losses, PRG, coil driver noise etc) can mean no observable squeezing. 
  • Actually, this entire discussion is moot unless we can get the RIN of the light incident on the PRM lower than the current level (estimated from MC2 transmission, filtered by CARM pole and ARM zero) by a factor of 60dB.
    • This is because even if we have 1mW contrast defect light leaking through the OMC, the beating of this field (in the amplitude quadrature) with the 20mW LO RIN (also almost entirely in the amplitude quad) yields significant noise contribution at 100 Hz (see Attachment #1).
    • Actually, we could have much more contrast defect leakage, as we have not accounted for asymmetries like arm loss imbalance.
    • So we need an ISS that has 60dB of gain at 100 Hz. 
    • The requirement on LO RIN is consistent with Eq 12 of this paper.
  • There is probably room to optimize SRC detuning and homodyne angle for each of these scenarios - for now, we just took the optimized combo for scenario #1 for evaluating all four scenarios.
  • OMC displacement noise seems to only be at the level of 1e-22 m/rtHz, assuming that the detuning for s-pol and p-pol is ~30 kHz if we were to lock at the middle of the two resonances
    • This assumes 0.02 deg difference in amplitude reflectivity b/w polarizations per optic, other parameters taken from aLIGO OMC design numbers.
    • We took OMC displacement noise from here.

Main unbudgeted noises:

  • Scattered light.
  • Angular control noise reinjection (not sure about the RP angular dynamics for the higher power yet).
  • Shot noise due to vacuum leaking from sym port (= DC contrast defect), but we expect this to not be significant at the level of the other noises in Atm #1.
  • Osc amp / phase.
  • AUX DoF cross coupling into DARM readout.
  • Laser frequency noise (although we should be immune to this because of our homodyne angle choice).

Threat matrix has been updated.

Attachment 1: PonderSqueeze_NB_LORIN.pdf
PonderSqueeze_NB_LORIN.pdf
  14224   Tue Oct 2 18:50:53 2018 yukiConfigurationASCPZT mirror calibration

[ Yuki, Gautam ]

I calibrated PZT mirrors. The ROUGH result was attached. (Note that some errors and trivial couplings coming from inclination of QPD were not considered here. This should be revised and posted again.) 

The PZT mirrors I calibrated were:

  • A 2-inch CVI mirror (45 degree, HR and AR for 532nm)
  • A 1-inch Laseroptik mirror (45 degree, HR and AR for 532nm)

I did the calibration as follows:

  • +-15V was supplied to PZT driver circuit, +100V to PZT driver bias, and +-18V to QPD amplifier.
  • Optical path length was set to be same as that when I calibrated QPD, which is 36cm.
  • The full range of CVI mirror is 3.5mrad according to its datasheet and linear range of QPD is 0.2mm, so I set the distance between PZT mirrors and QPD to be about 6cm. (I realized it was wrong. When mirror tilts 1 deg, the angle of beam changes 2 deg. So the distance should be the half.)
  • After applying 0V to PZT driver input (at that time 50V was applied to PZT mirror), then centered beam spot on QPD using steering mirror, which was confirmed by monitored Pitch and Yaw signals of QPD that were around zero.  
  • In order to avoid hysteresis effect, I stared with an input signal of -10V. I then increased the input voltage in steps of 1V through the full range from -10V to +10V DC. The other input was kept 0V.
  • Both the X and Y coordinates were noted in the plot in order to investigate pitch-yaw coupling.

The calibration factor was

CVI-pitch: 0.089 mrad/V

CVI-yaw: 0.096 mrad/V

Laseroptic-pitch: 0.062 mrad/V

Laseroptic-yaw: 0.070 mrad/V

Comments:

  • I made sure that PZT mirrors move linearly in full input range (+-10V).
  • PZT CH1 input: Yaw, CH2: Pitch, CH3: +100V bias
  • The calibration factor of PZT mirrors [mrad/V] are not consistent with previous calibration (elog:40m/8967). I will check it again.
  • I measured the beam power in order to calibrate QPD responce with a powermeter, but it didn't have high precision. So I used SUM output of QPD to the calibration.
  • Full range of PZT mirrors looks 2 times smaller.

Reference:
Previous calibration of the same mirrors, elog:40/8967

Attachment 1: PZTM1calibrationCH2.pdf
PZTM1calibrationCH2.pdf
Attachment 2: PZTM1calibrationCH1.pdf
PZTM1calibrationCH1.pdf
Attachment 3: PZTM2calibrationCH2.pdf
PZTM2calibrationCH2.pdf
Attachment 4: PZTM2calibrationCH1.pdf
PZTM2calibrationCH1.pdf
  14223   Mon Oct 1 22:20:42 2018 gautamUpdateSUSPrototyping HV Bias Circuit

Summary:

I've been plugging away at Altium prototyping the high-voltage bias idea, this is meant to be a progress update.

Details:

I need to get footprints for some of the more uncommon parts (e.g. PA95) from Rich before actually laying this out on a PCB, but in the meantime, I'd like feedback on (but not restricted to) the following:

  1. The top-level diagram: this is meant to show how all this fits into the coil driver electronics chain.
    • The way I'm imagining it now, this (2U) chassis will perform the summing of the fast coil driver output to the slow bias signal using some Dsub connectors (existing slow path series resistance would simply be removed). 
    • The overall output connector (DB15) will go to the breakout board which sums in the bias voltage for the OSEM PDs and then to the satellite box.
    • The obvious flaw in summing in the two paths using a piece of conducting PCB track is that if the coil itself gets disconnected (e.g. we disconnect cable at the vacuum flange), then the full HV appears at TP3 (see pg2 of schematic). This gets divided down by the ratio of the series resistance in the fast path to slow path, but there is still the possibility of damaging the fast-path electronics. I don't know of an elegant design to protect against this.
  2. Ground loops: I asked Johannes about the Acromag DACs, and apparently they are single ended. Hopefully, because the Sorensens power Acromags, and also the eurocrates, we won't have any problems with ground loops between this unit and the fast path.
  3. High-voltage precautons: I think I've taken the necessary precautions in protecting against HV damage to the components / interfaced electronics using dual-diodes and TVSs, but someone more knowledgable should check this. Furthermore, I wonder if a Molex connector is the best way to bring in the +/- HV supply onto the board. I'd have liked to use an SHV connector but can't find a comaptible board-mountable connector.
  4.  Choice of HV OpAmp: I've chosen to stick with the PA95, but I think the PA91 has the same footprint so this shouldn't be a big deal.
  5.  Power regulation: I've adapted the power regulation scheme Rich used in D1600122 - note that the HV supply voltage doesn't undergo any regulation on the board, though there are decoupling caps close to the power pins of the PA95. Since the PA95 is inside a feedback loop, the PSRR should not be an issue, but I'll confirm with LTspice model anyways just in case.
  6. Cost: 
    • ​​Each of the metal film resistors that Rich recommended costs ~$15.
    • The voltage rating on these demand that we have 6 per channel, and if this works well, we need to make this board for 4 optics.
    • The PA95 is ~$150 each, and presumably the high voltage handling resistors and capacitors won't be cheap.
    • Steve will update about his HV supply investigations (on a secure platform, NOT the elog), but it looks like even switching supplies cost north of $1200.
    • However, as I will detail in a separate elog, my modeling suggests that among the various technical noises I've modeled so far, coil driver noise is still the largest contribution which actually seems to exceed the unsqueezed shot noise of ~ 8e-19 m/rtHz for 1W input power and PRG 40 with 20ppm RT arm losses, by a smidge (~9e-19 m/rtHz, once we take into account the fast and slow path noises, and the fact that we are not exactly Johnson noise limited).

I also don't have a good idea of what the PCB layer structure (2 layers? 3 layers? or more?) should be for this kind of circuit, I'll try and get some input from Rich.

*Updated with current noise (Attachment #2) at the output for this topology of series resistance of 25 kohm in this path. Modeling was done (in LTspice) with a noiseless 25kohm resistor, and then I included the Johnson noise contribution of the 25k in quadrature. For this choice, we are below 1pA/rtHz from this path in the band we care about. I've also tried to estimate (Attachment #3) the contribution due to (assumed flat in ASD) ripple in the HV power supply (i.e. voltage rails of the PA95) to the output current noise, seems totally negligible for any reasonable power supply spec I've seen, switching or linear.

Attachment 1: CoilDriverBias.pdf
CoilDriverBias.pdf CoilDriverBias.pdf CoilDriverBias.pdf
Attachment 2: currentNoise.pdf
currentNoise.pdf
Attachment 3: PSRR.pdf
PSRR.pdf
  14222   Mon Oct 1 20:39:09 2018 gautamConfigurationASCc1asy

We need to set up a copy of the c1asx model (which currently runs on c1iscex), to be named c1asy, on c1iscey for the green steering PZTs. The plan discussed at the meeting last Wednesday was to rename the existing model c1tst into c1asy, and recompile it with the relevant parts copied over from c1asx. However, I suspect this will create some problems related to the "dcuid" field in the CDS params block (I ran into this issue when I tried to use the dcuid for an old model which no longer exists, called c1imc, for the c1omc model).

From what I can gather, we should be able to circumvent this problem by deleting the .par file corresponding to the c1tst model living at /opt/rtcds/caltech/c1/target/gds/param/, and rename the model to c1asy, and recompile it. But I thought I should post this here checking if anyone knows of other potential conflicts that will need to be managed before I start poking around and breaking things. Alternatively, there are plenty of cores available on c1iscey, so we could just set up a fresh c1asy model...

 
  • (write programming code of making alignment control automatically)
  14221   Mon Oct 1 13:33:55 2018 yukiConfigurationASCQPD calibration
Quote:

I assume this QPD set is a D1600079/D1600273 combo.

How much was the SUM output during the measurement? Also how much were the beam radii of this beam (from the error func fittings)?
Then the calibration [V/m] is going to be the linear/inv-linear function of the incident power and the beam radus.

You mean the linear range is +/-50mV (for a given beam), I guess.

  • The SUM output was from -174 to -127 mV.
  • The beam radii calculated from the error func fittings was 0.47 mm.
  • Total optical path length measured by a ruler= 36 cm.
  • Beam power measured at QPD was 2.96 mW. (There are some loss mechanism in the setup.)

Then the calibration factor of the QPD is

X axis: 584 * (POWER / 2.96mW) * (0.472mm /  RADIUS) [mV/mm]
Y axis: 588 * (POWER / 2.96mW) * (0.472mm /  RADIUS) [mV/mm].

Attachment 1: Pic_QPDcalibration.jpg
Pic_QPDcalibration.jpg
  14220   Mon Oct 1 12:03:41 2018 not yukiConfigurationASCPZT driver board verification

I assume this QPD set is a D1600079/D1600273 combo.

How much was the SUM output during the measurement? Also how much were the beam radii of this beam (from the error func fittings)?
Then the calibration [V/m] is going to be the linear/inv-linear function of the incident power and the beam radus.

You mean the linear range is +/-50mV (for a given beam), I guess.

 

  14219   Sun Sep 30 20:14:51 2018 yukiConfigurationASCQPD calibration

[ Yuki, Gautam, Steve ]

Results:
I calibrated a QPD (D1600079, V1009) and made sure it performes well. The calibration constants are as follows:

X-Axis: 584 mV/mm
Y-Axis: 588 mV/mm

Details:
The calibration of QPD is needed to calibrate steeing PZT mirrors. It was measured by moving QPD on a translation stage. The QPD was connected to its amplifier (D1700110-v1) and +-18V was supplied from DC power supplier. The amplifier has three output ports; Pitch, Yaw, and Sum. I did the calibration as follows:

  • Center beam spot on QPD using steering mirror, which was confirmed by monitored Pitch and Yaw signals that were around zero.  
  • Kept Y-axis micrometer fixed, moved X-axis micrometer and measured the outputs. 
  • Repeated the procedure for the Y-axis. 

The results are attached. The main signal was fitted with error function and I drawed a slope at zero crossing point, which is calibration factor. I determined the linear range of the QPD to be when the output was in range -50V to 50V, then corresponding displacement range is about 0.2 mm width. Using this result, the PZT mirrors will be calibrated in linear range of the QPD tomorrow. 

Comments:

  • Some X-Y coupling existed. When one axis micrometer was moved, a little signal of the other direction was also generated.
  • As Gautam proposed in the previous study, there is some hysteresis. That process would bring some errors to this result.
  • A scale of micrometer is expressed in INCH!
  • The micrometer I used was made to have 1/2 inch range, but it didn't work well and the range of X-axis was much narrower. 

Reference:
previous experiment by Gautam for X-arm: elog:40m/8873, elog:40m/8884

Attachment 1: QPDcalibrationXaxis.pdf
QPDcalibrationXaxis.pdf
Attachment 2: QPDcalibrationYaxis.pdf
QPDcalibrationYaxis.pdf
  14218   Thu Sep 27 14:02:55 2018 yukiConfigurationASCPZT driver board verification

[ Yuki, Gautam ]

I fixed the input terminal that had been off, and made sure PZT driver board performs as we expect. 

At first I ran a simulation of the PZT driver circuit using LTspice (Attached #1 and #2). It shows that when the bias is 30V the driver performs well only with high input volatage (bigger than 3V). Then I measured the performance as following way:

  1. Applied +-15V to the board with an expansion card and 31.8V to the high voltage port which is the maximum voltage of PS280 DC power supplier C10013.
  2. Terminated input and connectd input bias to GND, then set offset to -10.4V. This value is refered as elog:40m/8832.
  3. Injected DC signal into input port using a function generator.
  4. Measured voltage at the OUT port and MON port.

The result of this is attached #3 and #4. It is consistent with simulated one. All ports performed well.

  • V(M1_PIT_OUT) = -4.86 *Vin +49.3 [V]
  • V(M1_YAW_OUT) = -4.86 *Vin +49.2 [V]
  • V(M2_PIT_OUT) = -4.85 *Vin +49.4 [V]
  • V(M2_YAW_OUT) = -4.86 *Vin +49.1 [V]
  • V(M1_PIT_MON) = -0.333 *Vin +3.40 [V]
  • V(M1_YAW_MON) = -0.333 *Vin +3.40 [V]
  • V(M2_PIT_MON) = -0.333 *Vin +3.40 [V]
  • V(M2_YAW_MON) = -0.333 *Vin +3.40 [V]

The high voltage points (100V DC) remain to be tested.

Attachment 1: PZTdriverSimulationDiagram.pdf
PZTdriverSimulationDiagram.pdf
Attachment 2: PZTdriverSimulationResult.pdf
PZTdriverSimulationResult.pdf
Attachment 3: PZTdriverPerformanceCheck_ResultOUT.pdf
PZTdriverPerformanceCheck_ResultOUT.pdf
Attachment 4: PZTdriverPerformanceCheck_ResultMON.pdf
PZTdriverPerformanceCheck_ResultMON.pdf
Attachment 5: PZTdriver.asc
Version 4
SHEET 1 2120 2120
WIRE 1408 656 1408 624
WIRE 1552 656 1552 624
WIRE 1712 656 1712 624
WIRE 1872 656 1872 624
WIRE 2016 656 2016 624
WIRE 1408 768 1408 736
WIRE 1552 768 1552 736
WIRE 1712 768 1712 736
... 193 more lines ...
  14217   Wed Sep 26 10:07:16 2018 SteveUpdateVACwhy reboot c1vac1

Precondition: c1vac1 & c1vac2 all LED warning lights green [ atm3 ], the only error message is in the gauge readings NO COMM, dataviewer will plot zero [ atm1 ], valves are operational

When our vacuum gauges read " NO COMM " than our INTERLOCKS  do  NOT communicate either.

So V1 gate valve and PSL output shutter can not be triggered to close if the the IFO pressure goes up.                        

   [ only CC1_HORNET_PRESSURE reading is working in this condition because it goes to a different compuer ] 

Quote:

[steve, gautam]

Following the procedure in this elog, we effected a reset of the vacuum slow machines. Usually, I just turn the key on these crates to do a power cycle, but Steve pointed out that for the vacuum machines, we should only push the "reset" button.

While TP1 was spun down, we took the opportunity to replace the TP1 controller with a spare unit the company has sent us for use while our unit is sent to them for maintenance. The procedure was in principle simple (I only list the additional ones, for the various valve closures, see the slow machine reset procedure elog):

  • Turn power off using switch on rear.
  • Remove 4 connecting cables on the back.
  • Switch controllers.
  • Reconnect 4 cables on the back panel.
  • Turn power back on using switch on rear.

However, we were foiled by a Philips screw on the DB37 connector labelled "MAG BRG", which had all its head worn out. We had to make a cut in this screw using a saw blade, and use a "-" screwdriver to get this troublesome screw out. Steve suspects this is a metric gauge screw, and will request the company to send us a new one, we will replace it when re-installing the maintaiend controller. 

Attachments #1 and #2 show the Vacuum MEDM screen before and after the reboot respectively - evidently, the fields that were reading "NO COMM" now read numbers. Attachment #3 shows the main volume pressure during this work.

Quote:

The problem will be revisited on Monday.

 

Attachment 1: NOcomm.png
NOcomm.png
Attachment 2: Reboot_&_sawp.png
Reboot_&_sawp.png
Attachment 3: c1vac1&2_.jpg
c1vac1&2_.jpg
  14216   Tue Sep 25 18:08:50 2018 yukiConfigurationASCY end table upgrade plan

[ Yuki, Gautam ]

We want to remotely control steeing PZT mirrors so its driver is needed. We already have a PZT driver board (D980323-C) and the output voltage is expected to be verified to be in the range 0-100 V DC for input voltages in the range -10 to 10 V DC.
Then I checked to make sure ir perform as we expected. The input signal was supplied using voltage calibrator and the output was monitored using a multimeter. 
But it didn't perform well. Some tuning of voltage bias seemed to be needed. I will calculate its transfer function by simulation and check the performance again tommorow. And I found one solder was off so it needs fixing.  

Reference:
diagram --> elog 8932
 

Plan of Action:

  • Check PZT driver performs as we expected
  • Also check cable, high voltage, PZT mirrors, anti-imaging board
  • Obtain calibration factor of PZT mirrors using QPD
  • Measure some status value before changing setup (such as tranmitted power of green laser)
  • Revise setup after a new lens arrives
  • Align the setup and check mode-matching
  • Measure status value again and confirm it improves
  • (write programming code of making alignment control automatically)
  14215   Mon Sep 24 15:06:10 2018 gautamUpdateVACc1vac1 reboot + TP1 controller replacement

[steve, gautam]

Following the procedure in this elog, we effected a reset of the vacuum slow machines. Usually, I just turn the key on these crates to do a power cycle, but Steve pointed out that for the vacuum machines, we should only push the "reset" button.

While TP1 was spun down, we took the opportunity to replace the TP1 controller with a spare unit the company has sent us for use while our unit is sent to them for maintenance. The procedure was in principle simple (I only list the additional ones, for the various valve closures, see the slow machine reset procedure elog):

  • Turn power off using switch on rear.
  • Remove 4 connecting cables on the back.
  • Switch controllers.
  • Reconnect 4 cables on the back panel.
  • Turn power back on using switch on rear.

However, we were foiled by a Philips screw on the DB37 connector labelled "MAG BRG", which had all its head worn out. We had to make a cut in this screw using a saw blade, and use a "-" screwdriver to get this troublesome screw out. Steve suspects this is a metric gauge screw, and will request the company to send us a new one, we will replace it when re-installing the maintaiend controller. 

Attachments #1 and #2 show the Vacuum MEDM screen before and after the reboot respectively - evidently, the fields that were reading "NO COMM" now read numbers. Attachment #3 shows the main volume pressure during this work.

Quote:

The problem will be revisited on Monday.

Attachment 1: beforeReboot.png
beforeReboot.png
Attachment 2: afterReboot.png
afterReboot.png
Attachment 3: CC1.png
CC1.png
  14214   Mon Sep 24 11:09:05 2018 yukiConfigurationASCY end table upgrade plan

[ Yuki, Steve ]

With Steve's help, we checked a new lens can be set soon after dichroic mirror.

Quote:

There may be a problem: One lens should be put soon after dichroic mirror, but there is little room for fix it. (Attached #4, It will be put where the pedestal is.)  Tomorrow we will check this problem again.

Attachment 1: pic0924_1.jpg
pic0924_1.jpg
  14213   Sun Sep 23 20:15:35 2018 KojiSummaryOMCMontecarlo simulation of the phase difference between P and S pols for a modeled HR mirror

Link to OMC_Lab ELOG 308

  14212   Sun Sep 23 19:32:23 2018 yukiConfigurationASCY end table upgrade plan

[ Yuki, Gautam ]

The setup I designed before has abrupt gouy phase shift between two steering mirrors which makes alignment much sensitive. So I designed a new one (Attached #1, #2 and #3). It improves the slope of gouy phase and the difference between steering mirrors is about 100 deg. To install this, we need new lenses: f=100mm, f=200mm, f=-250mm which have 532nm coating. If this setup is OK, I will order them.

There may be a problem: One lens should be put soon after dichroic mirror, but there is little room for fix it. (Attached #4, It will be put where the pedestal is.)  Tomorrow we will check this problem again.

And another problem; one steering mirror on the corner of the box is not easy to access. (Attached #5) I have to design a new seup with considering this problem.

Quote:

One of the example for improvement is just adding a new lens (f=10cm) soon after the doubling crystal. That will make mode matching better (100%) and also make separation better (85 deg) (Attachments #4 and #5). I'm checking whether we have the lens and there is space to set it. And I will measure current power of transmitted main laser in order to confirm the improvement of alignment.

 

Attachment 1: Pic_NewSetup0923_AUXYgreen.jpeg
Pic_NewSetup0923_AUXYgreen.jpeg
Attachment 2: ModeMatchingSolution_Result.pdf
ModeMatchingSolution_Result.pdf
Attachment 3: ModeMatchingSolution_Magnified_0923.jpg
ModeMatchingSolution_Magnified_0923.jpg
Attachment 4: pic0923_1.jpg
pic0923_1.jpg
Attachment 5: pic0923_2.jpg
pic0923_2.jpg
  14211   Sun Sep 23 17:38:48 2018 yukiUpdateASCAlignment of AUX Y end green beam was recovered

[ Yuki, Koji, Gautam ]

An alignment of AUX Y end green beam was bad. With Koji and Gautam's advice, it was recovered on Friday. The maximum value of TRY was about 0.5.

  14210   Sat Sep 22 00:21:07 2018 KojiUpdateCDSFrequent time out

[Gautam, Koji]

We had another crash of c1sus and Gautam did full power cycling of c1sus. It was a sturggle to recover all the frontends, but this solved the timing issue.

We went through full reset of c1sus, and rebooting all the other RT hosts, as well as daqd and fb1.

Attachment 1: 23.png
23.png
  14208   Fri Sep 21 19:50:17 2018 KojiUpdateCDSFrequent time out

Multiple realtime processes on c1sus are suffering from frequent time outs. It eventually knocks out c1sus (process).

Obviously this has started since the fiber swap this afternoon.

gautam 10pm: there are no clues as to the origin of this problem on the c1sus frontend dmesg logs. The only clue (see Attachment #3) is that the "ADC" error bit in the CDS status word is red - but opening up the individual ADC error log MEDM screens show no errors or overflows. Not sure what to make of this. The IOP model on this machine (c1x02) reports an error in the "Timing" bit of the CDS status word, but from the previous exchange with Rolf / J Hanks, this is down to a misuse of ADC0 Ch31 which is supposed to be reserved for a DuoTone diagnostic signal, but which we use for some other signal (one of the MC suspension shadow sensors iirc). The response is also not consistent with this CDS manual - which suggests that an "ADC" error should just kill the models. There are no obvious red indicator lights in the c1sus expansion chassis either.

Attachment 1: 33.png
33.png
Attachment 2: 49.png
49.png
Attachment 3: Screenshot_from_2018-09-21_21-52-54.png
Screenshot_from_2018-09-21_21-52-54.png
  14207   Fri Sep 21 16:51:43 2018 gautamUpdateVACc1vac1 is unresponsive

Steve pointed out that some of the vacuum MEDM screen fields were reporting "NO COMM". Koji confirmed that this is a c1vac1 problem, likely the same as reported here and can be fixed using the same procedure.

However, Steve is worried that the interlock won't kick in in case of a vacuum emergency, so we are leaving the PSL shutter closed over the weekend. The problem will be revisited on Monday.

  14206   Fri Sep 21 16:46:38 2018 gautamUpdateCDSNew PCIe fiber installed and routed

[steve, koji, gautam]

We took another pass at this today, and it seems to have worked - see Attachment #1. I'm leaving CDS in this configuration so that we can investigate stability. IMC could be locked. However, due to the vacuum slow machine having failed, we are going to leave the PSL shutter closed over the weekend.

Attachment 1: PCIeFiber.png
PCIeFiber.png
Attachment 2: IMG_5878.JPG
IMG_5878.JPG
  14205   Fri Sep 21 09:59:09 2018 yukiConfigurationASCY end table upgrade plan

[Yuki, Gautam]

Attachments #1 is the current setup of AUX Y Green locking and it has to be improved because:

  • current efficiency of mode matching is about 50%
  • current setup doesn't separate the degrees of freedom of TEM01 with PZT mirrors (the difference of gouy phase between PZT mirrors should be around 90 deg) 
  • we want to remotely control PZT mirrors for alignment
    (Attachments #2 and #3)

About the above two: 

One of the example for improvement is just adding a new lens (f=10cm) soon after the doubling crystal. That will make mode matching better (100%) and also make separation better (85 deg) (Attachments #4 and #5). I'm checking whether we have the lens and there is space to set it. And I will measure current power of transmitted main laser in order to confirm the improvement of alignment.

About the last:

I am considering what component is needed. 

Reference:

Attachment 1: Pic_CurrentSetup_AUXYgreen.jpeg
Pic_CurrentSetup_AUXYgreen.jpeg
Attachment 2: ModeMatchingSolution_Current.pdf
ModeMatchingSolution_Current.pdf
Attachment 3: ModeMatchingSolution_Current_Magnified.pdf
ModeMatchingSolution_Current_Magnified.pdf
Attachment 4: ModeMatchingSolution_Optimized.pdf
ModeMatchingSolution_Optimized.pdf
Attachment 5: ModeMatchingSolution_Optimized_Magnified.pdf
ModeMatchingSolution_Optimized_Magnified.pdf
  14203   Thu Sep 20 16:19:04 2018 gautamUpdateCDSNew PCIe fiber install postponed to tomorrow

[steve, gautam]

This didn't go as smoothly as planned. While there were no issues with the new fiber over the ~3 hours that I left it plugged in, I didn't realize the fiber has distinct ends for the "HOST" and "TARGET" (-5 points to me I guess). So while we had plugged in the ends correctly (by accident) for the pre-lunch test, while routing the fiber on the overhead cable tray, we switched the ends (because the "HOST" end of the cable is close to the reel and we felt it would be easier to do the routing the other way. 

Anyway, we will fix this tomorrow. For now, the old fiber was re-connected, and the models are running. IMC is locked.

Quote:

Pictures + more procedural details + proper routing of the protected fiber along cable trays after lunch. If this doesn't help the stability problem, we are out of ideas again, so fingers crossed...

  14202   Thu Sep 20 11:29:04 2018 gautamUpdateCDSNew PCIe fiber housed

[steve, yuki, gautam]

The plastic tubing/housing for the fiber arrived a couple of days ago. We routed ~40m of fiber through roughly that length of the tubing this morning, using some custom implements Steve sourced. To make sure we didn't damage the fiber during this process, I'm now testing the vertex models with the plastic tubing just routed casually (= illegally) along the floor from 1X4 to 1Y3 (NOTE THAT THE WIKI PAGE DIAGRAM IS OUT OF DATE AND NEEDS TO BE UPDATED), and have plugged in the new fiber to the expansion chassis and the c1lsc front end machine. But I'm seeing a DC error (0x4000), which is indicative of some sort of timing error (Attachment #1) **. Needs more investigation...

Pictures + more procedural details + proper routing of the protected fiber along cable trays after lunch. If this doesn't help the stability problem, we are out of ideas again, so fingers crossed...

** In the past, I have been able to fix the 0x4000 error by manually rebooting fb (simply restarting the daqd processes on fb using sudo systemctl restart daqd_* doesn't seem to fix the problem). Sure enough, seems to have done the job this time as well (Attachment #2). So my initial impression is that the new fiber is functioning alright yes.

Quote:

The PCIe fiber replacement is a more involved project (Steve is acquiring some protective tubing to route it from the FE in 1X6 to the expansion chassis in 1Y3)

Attachment 1: PCIeFiberSwap.png
PCIeFiberSwap.png
Attachment 2: PCIeFiberSwap_FBrebooted.png
PCIeFiberSwap_FBrebooted.png
  14201   Thu Sep 20 08:17:14 2018 SteveUpdateSUSlocal 3.4M earth quake

M3.4 Colton shake did not trip sus.

 

Attachment 1: local_3.4M.png
local_3.4M.png
  14200   Tue Sep 18 17:56:01 2018 not gautamUpdateIOOPMC and IMC relocked, WFS inputs turned off

I restarted the LSC models in the usual way via the c1lsc reboot script. After doing this I was able to lock the YARM configuration for more noise coupling scripting.

Quote:

The PMC and IMC were unlocked. Both were re-locked, and alignment of both cavities were adjusted so as to maximize MC2 trans (by hand, input alignment to PMC tweaked on PSL table, IMC alignment tweaked using slow bias voltages). I disabled the inputs to the WFS loops, as it looks like they are not able to deal with the glitching IMC suspensions. c1lsc models have crashed again but I am not worrying about that for now.

9pm: The alignment is wandering all over the place so I'm just closing the PSL shutter for now.

 

  14199   Tue Sep 18 14:02:37 2018 SteveUpdatesafety safety training

Yuki Miyazaki received 40m specific basic safety training.

 

  14198   Mon Sep 17 12:28:19 2018 gautamUpdateIOOPMC and IMC relocked, WFS inputs turned off

The PMC and IMC were unlocked. Both were re-locked, and alignment of both cavities were adjusted so as to maximize MC2 trans (by hand, input alignment to PMC tweaked on PSL table, IMC alignment tweaked using slow bias voltages). I disabled the inputs to the WFS loops, as it looks like they are not able to deal with the glitching IMC suspensions. c1lsc models have crashed again but I am not worrying about that for now.

9pm: The alignment is wandering all over the place so I'm just closing the PSL shutter for now.

  14197   Wed Sep 12 22:22:30 2018 KojiUpdateComputersSSL2.0, SSL3.0 disabled

LIGO GC notified us that nodus had SSL2.0 and SSL3.0 enabled. This has been disabled now.
The details are described on 40m wiki.

  14196   Mon Sep 10 12:44:48 2018 JonUpdateCDSADC replacement in c1lsc expansion chassis

Gautam and I restarted the models on c1lsc, c1ioo, and c1sus. The LSC system is functioning again. We found that only restarting c1lsc as Rolf had recommended did actually kill the models running on the other two machines. We simply reverted the rebootC1LSC.sh script to its previous form, since that does work. I'll keep using that as required until the ongoing investigations find the source of the problem.

Quote:

Looks like the ADC was not to blame, same symptoms persist.

Quote:

The PCIe fiber replacement is a more involved project (Steve is acquiring some protective tubing to route it from the FE in 1X6 to the expansion chassis in 1Y3), but hopefully the problem was the ADC card with red indicator light, and replacing it has solved the issue.

 

ELOG V3.1.3-